OpenStudy (anonymous):

factor the sum and differences of cubes. 8/27 + m^6

OpenStudy (anonymous):

I am not sure of what you want but this is the factorization of your expression: \[ \frac{8}{27} + m^6 = \frac{1}{27} \left(3 m^2+2\right) \left(9 m^4 -6 m^2+4\right) \]

OpenStudy (anonymous):

this is great ty again

OpenStudy (anonymous):

Glad to help :)